# Enhancing LLM Reliability via Explicit Knowledge Boundary Modeling

### v2

@@ Appendix A Refinement Model
Appendix B Baselines Details
@@
Token Probability (Prob): Computes average token-level probabilities, classifying outputs as "sure" or "unsure" based on a threshold , with probabilities normalized and set to 0.8.
Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as "sure," while others are marked as "unsure." We conduct ten samples () and set to 50%.
P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding "True" than "False" as "sure," otherwise as "unsure."
Appendix C Scalability Analysis
@@
| Give you a dialog history between USER and SYSTEM, and a slot-value pair extracted from the dialog, I want you to analyze on the dialog history and review/refine the given dialog state slot-value pair to make them more accurate and reliable. ## Domain Slot Spaces { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Question Dialog history: USER: Hello. I need train to London liverpool Street. SYSTEM: Where are you departing from and do you have a time preference? USER: Yes, I’d like to leave Cambridge sometime after 7:00. SYSTEM: I have one that leaves at 7:59 and 4 more that depart every two hours after. USER: The 7:59 will be fine. Predicted slot-value pair: train: leaveat = 7:59 Refinement: Thinking steps: step1: structure aspect: "leaveat" is one of the slots in the train domain and the value is incorrectly specified as "7:59" (formation error). Precise time should strictly follow the "HH:MM" format. step2: semantic aspect: "leaveat" is not a "name"-related slot, so the value must be extracted from the user’s direct intent. Although the system provides a train that departs at 7:59, the user only mentions that he’d like to leave Cambridge after 7:00. Considering the formation, the correct value should be 07:00. step3: domain train correct; slot leaveat correct; value 7:59 incorrect. step4: the predicted slot-value pair is incorrect in both structure and semantic, and the value should be corrected. Refine action: correct Refined slot-value pair: train: leaveat = 07:00 |
@@
| Give you a dialog history between USER and SYSTEM, a predicted slot-value pair and a gold refinement, I want you to analyze on the dialog history, refer to the examples I’ve provided and the golden refinement, and then continue to provide the thinking steps following the predefined principles. Firstly, we use json dict to describe the slots and their corresponding value space. Then, we will specify the requirements you need to comply and provide some examples. Last, we will present you with the dialog history, the predicted dialog state slot-value pair and the gold refinement for you to perform continuation. ## Domain Slot Spaces { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Meta Requirements 1. Analyze the dialog history and the predicted slot-value pair carefully, correct and refine the pair according to the following Detailed Refinement Principles. 2. For the given slot-value pair, you should only take one of the actions of "reserve", "remove" or "correct" according to the following principles. a. reserve: if the slot is relevant to the dialog history and the value is correctly specified, you should keep the pair. b. remove: if the slot is irrelevant to the given dialog history or there’s not corresponding value could be extracted from the history, you should print "none" indicating that the pair has been removed. c. correct: if the slot is relevant to the dialog history and the value is incorrectly specified, you should extract the correct value and replace it. 3. You need to think and reason in the same way as the Examples. Please refer to the Examples for formatting requirement. Make sure the output is all lowercase. Only focus on the given slot-value pair and do not consider any other slot. only provide the reasonable thinking steps, do not provide any extra prefixes or suffixes. ## Detailed Refinement Principles ## Examples ## Question Dialog history: USER: … SYSTEM: … Predicted slot-value pair: train: leaveat = 7:59 Refinement: Refine action: correct Refined slot-value pair: train: leaveat = 07:00 Thinking steps: |
Appendix D Refinement Cost
The detailed overall results of cost are presented in Table 3. The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 7 and Figure 8, respectively.
Appendix E LLM Instructions
@@
E.1 Inference Instruction
The prompts used with the original untrained LLAMA3 model are shown in Tables 7 and 8. Among them, the Direct Inference Instruction requires the model to directly output the prediction results, while the Reliability Inference Instruction additionally requires the model to output a confidence label alongside the prediction.
@@ E.1 Inference Instruction
| Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state. Firstly, we use json dict to describe the slots and their corresponding value space in each domain. Then, we will specify the requirements you need to comply. Last, we demonstrate some use cases. ## Domain and Slot Space The availabel ontology of the dialog state is as follows: { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Requirements 1. Analyze the dialog history carefully and fill the relevant domains and slots. 2. For slots with a specified value range, responses must fall within the provided range. For slots without specified value range, the answer must be extracted from the history. Set value as "dontcare" if user doesn’t have a preference. 3. You only need to consider the domains and slots that are relevant to the conversation history. Do not include those irrelevant in your response and avoid presenting empty domains or slots. 4. Your answer should also be in one-line jsonl format and make sure the output is all lowercase. Do not provide any extra prefixes or suffixes or any explanations. ## Output Dialog State Example {"hotel": {"area": "centre", "name": "alexander bed and breakfast", "parking": "yes", "type": "guesthouse"}, "attraction": {"name": "kambar"}} ## Examples shots… |
@@
| Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state. Firstly, we use json dict to describe the slots and their corresponding value space in each domain. Then, we will specify the requirements you need to comply. Last, we demonstrate some use cases. ## Domain and Slot Space The availabel ontology of the dialog state is as follows: { "hotel": { "name": { "data_type": str, "example": "hamilton lodge" }, … }, … } ## Requirements 1. Analyze the dialog history carefully and fill the relevant domains and slots. 2. For slots with a specified value range, responses must fall within the provided range. For slots without specified value range, the answer must be extracted from the history. Set value as "dontcare" if user doesn’t have a preference. 3. You should try to cover as many domians and slot-value pairs relevant to the conversation history as possible. Mark each slot with either "sure" or "unsure" according to your confidence in it. For slot-value pairs that you are pretty sure that they are truly relevant and should be included, and meanwhile you have great confidence in the correctness of the value, you should tag it with "sure". Otherwise, you should tag the slot-value pair with "unsure". 4. Principles: a. Goal one: Achieving as close to 100% accuracy as possible in those "sure" slot-value pairs. b. Goal two: The "sure" part along with the "unsure" part should cover all possible slots involved in the dialog history as much as possible. c. Heavy penalization: Providing incorrect slot-value pairs in the "sure" part. d. Heavy penalization: Missing slot-value pairs that should be extracted. e. Light penalization: Providing incorrect or redundant slot-value pairs in the "unsure" part. 5. Your answer should also be in one-line jsonl format and make sure the output is all lowercase. Do not provide any extra prefixes or suffixes or any explanations. Please refer to the Output Dialog State Example for formatting requirement. ## Output Dialog State Example {"hotel": {"area": {"value": "centre", "confidence": "sure"}, "name": {"value": "alexander bed and breakfast", "confidence": "sure"}, "parking": {"value": "yes", "confidence": "unsure"}, "type": {"value": "guesthouse", "confidence": "sure"}}, "attraction": {"name": {"value": "kambar", "confidence": "unsure"}}} ## Examples shots… |
E.2 Training Instruction
The prompts used with the trained LLAMA3 model are shown in Tables 9 and 10. The Direct Training Instruction is designed to train foundation models capable of directly outputting prediction results. In contrast, the Reliability Training Instruction is designed to train models capable of explicitly outputting confidence labels alongside predictions.

### v3

@@ Appendix C Refinement Model
Appendix D Baselines Details
@@
Token Probability (Prob): Computes average token-level probabilities, classifying outputs as ”sure” or ”unsure” based on a threshold , with probabilities normalized and set to 0.8.
Self-Consistency (SC): Samples times, classifying predictions that appear consistently at proportion as ”sure,” while others are marked as ”unsure.” We conduct ten samples () and set to 50%.
P(True): Based on predictions from the Direct baseline, this method evaluates the correctness of each prediction individually, marking those with a higher probability of responding ”True” than ”False” as ”sure,” otherwise as ”unsure.”
Appendix E Scalability Analysis
@@
Appendix F Refinement Cost
The detailed overall results of cost are presented in Table 7 and 7. The relationships between cost and performance for the BiTOD and SGD datasets are illustrated in Figure 8(a) and Figure 8(b), respectively.
Method Type Method MultiWOZ-2.4 BiTOD SGD Medium High Medium High Medium High Prompt Direct 100.00 100.00 100.00 100.00 100.00 100.00 Uncertainty Prob 6.55 5.74 3.29 1.37 8.14 5.53 SC 28.93 12.17 6.25 2.12 41.68 41.42 Reliability (Ours) SFT 13.52 5.12 2.16 0.66 12.60 1.87 + DPO Joint 10.70 3.94 1.36 1.04 11.66 2.67 + DPO Post 13.32 8.59 3.44 1.46 19.01 4.19
@@ Appendix F Refinement Cost
Give you a dialog history between USER and SYSTEM, and a slot-value pair extracted from the dialog, I want you to analyze on the dialog history and review/refine the given dialog state slot-value pair to make them more accurate and reliable. ## Domain Slot Spaces { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Question Dialog history: USER: Hello. I need train to London liverpool Street. SYSTEM: Where are you departing from and do you have a time preference? USER: Yes, I’d like to leave Cambridge sometime after 7:00. SYSTEM: I have one that leaves at 7:59 and 4 more that depart every two hours after. USER: The 7:59 will be fine. Predicted slot-value pair: train: leaveat = 7:59 Refinement: Thinking steps: step1: structure aspect: ”leaveat” is one of the slots in the train domain and the value is incorrectly specified as ”7:59” (formation error). Precise time should strictly follow the ”HH:MM” format. step2: semantic aspect: ”leaveat” is not a ”name”-related slot, so the value must be extracted from the user’s direct intent. Although the system provides a train that departs at 7:59, the user only mentions that he’d like to leave Cambridge after 7:00. Considering the formation, the correct value should be 07:00. step3: domain train correct; slot leaveat correct; value 7:59 incorrect. step4: the predicted slot-value pair is incorrect in both structure and semantic, and the value should be corrected. Refine action: correct Refined slot-value pair: train: leaveat = 07:00 |
@@
Give you a dialog history between USER and SYSTEM, a predicted slot-value pair and a gold refinement, I want you to analyze on the dialog history, refer to the examples I’ve provided and the golden refinement, and then continue to provide the thinking steps following the predefined principles. Firstly, we use json dict to describe the slots and their corresponding value space. Then, we will specify the requirements you need to comply and provide some examples. Last, we will present you with the dialog history, the predicted dialog state slot-value pair and the gold refinement for you to perform continuation. ## Domain Slot Spaces { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Meta Requirements 1. Analyze the dialog history and the predicted slot-value pair carefully, correct and refine the pair according to the following Detailed Refinement Principles. 2. For the given slot-value pair, you should only take one of the actions of ”reserve”, ”remove” or ”correct” according to the following principles. a. reserve: if the slot is relevant to the dialog history and the value is correctly specified, you should keep the pair. b. remove: if the slot is irrelevant to the given dialog history or there’s not corresponding value could be extracted from the history, you should print ”none” indicating that the pair has been removed. c. correct: if the slot is relevant to the dialog history and the value is incorrectly specified, you should extract the correct value and replace it. 3. You need to think and reason in the same way as the Examples. Please refer to the Examples for formatting requirement. Make sure the output is all lowercase. Only focus on the given slot-value pair and do not consider any other slot. only provide the reasonable thinking steps, do not provide any extra prefixes or suffixes. ## Detailed Refinement Principles ## Examples ## Question Dialog history: USER: … SYSTEM: … Predicted slot-value pair: train: leaveat = 7:59 Refinement: Refine action: correct Refined slot-value pair: train: leaveat = 07:00 Thinking steps: |
Appendix G LLM Instructions
@@
G.1 Inference Instruction
The prompts used with the original untrained LLAMA3 model are shown in Tables 10 and 11. Among them, the Direct Inference Instruction requires the model to directly output the prediction results, while the Reliability Inference Instruction additionally requires the model to output a confidence label alongside the prediction.
@@ G.1 Inference Instruction
Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state. Firstly, we use json dict to describe the slots and their corresponding value space in each domain. Then, we will specify the requirements you need to comply. Last, we demonstrate some use cases. ## Domain and Slot Space The availabel ontology of the dialog state is as follows: { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Requirements 1. Analyze the dialog history carefully and fill the relevant domains and slots. 2. For slots with a specified value range, responses must fall within the provided range. For slots without specified value range, the answer must be extracted from the history. Set value as ”dontcare” if user doesn’t have a preference. 3. You only need to consider the domains and slots that are relevant to the conversation history. Do not include those irrelevant in your response and avoid presenting empty domains or slots. 4. Your answer should also be in one-line jsonl format and make sure the output is all lowercase. Do not provide any extra prefixes or suffixes or any explanations. ## Output Dialog State Example {”hotel”: {”area”: ”centre”, ”name”: ”alexander bed and breakfast”, ”parking”: ”yes”, ”type”: ”guesthouse”}, ”attraction”: {”name”: ”kambar”}} ## Examples shots… |
@@
Give you a dialog history between USER and SYSTEM, I want you to analyze on it and generate the dialog state. Firstly, we use json dict to describe the slots and their corresponding value space in each domain. Then, we will specify the requirements you need to comply. Last, we demonstrate some use cases. ## Domain and Slot Space The availabel ontology of the dialog state is as follows: { ”hotel”: { ”name”: { ”data_type”: str, ”example”: ”hamilton lodge” }, … }, … } ## Requirements 1. Analyze the dialog history carefully and fill the relevant domains and slots. 2. For slots with a specified value range, responses must fall within the provided range. For slots without specified value range, the answer must be extracted from the history. Set value as ”dontcare” if user doesn’t have a preference. 3. You should try to cover as many domians and slot-value pairs relevant to the conversation history as possible. Mark each slot with either ”sure” or ”unsure” according to your confidence in it. For slot-value pairs that you are pretty sure that they are truly relevant and should be included, and meanwhile you have great confidence in the correctness of the value, you should tag it with ”sure”. Otherwise, you should tag the slot-value pair with ”unsure”. 4. Principles: a. Goal one: Achieving as close to 100% accuracy as possible in those ”sure” slot-value pairs. b. Goal two: The ”sure” part along with the ”unsure” part should cover all possible slots involved in the dialog history as much as possible. c. Heavy penalization: Providing incorrect slot-value pairs in the ”sure” part. d. Heavy penalization: Missing slot-value pairs that should be extracted. e. Light penalization: Providing incorrect or redundant slot-value pairs in the ”unsure” part. 5. Your answer should also be in one-line jsonl format and make sure the output is all lowercase. Do not provide any extra prefixes or suffixes or any explanations. Please refer to the Output Dialog State Example for formatting requirement. ## Output Dialog State Example {”hotel”: {”area”: {”value”: ”centre”, ”confidence”: ”sure”}, ”name”: {”value”: ”alexander bed and breakfast”, ”confidence”: ”sure”}, ”parking”: {”value”: ”yes”, ”confidence”: ”unsure”}, ”type”: {”value”: ”guesthouse”, ”confidence”: ”sure”}}, ”attraction”: {”name”: {”value”: ”kambar”, ”confidence”: ”unsure”}}} ## Examples shots… |
G.2 Training Instruction
The prompts used with the trained LLAMA3 model are shown in Tables 12 and 13. The Direct Training Instruction is designed to train foundation models capable of directly outputting prediction results. In contrast, the Reliability Training Instruction is designed to train models capable of explicitly outputting confidence labels alongside predictions.